How do I write out this problem to solve it.
Ellen is 11 years older than Maja. Last year Ellen was twice as old as Maja. How old is Maja now?

Answer by Alan3354(69443)

Ellen is 11 years older than Maja. Last year Ellen was twice as old as Maja. How old is Maja now?
--------------
E = M + 11
E-1 = 2*(M-1) --> E-1 = 2M-2
E = 2M - 1
-------------
M + 11 = 2M - 1
M = 12
-------
E = 23